"\iint" y/x^2+1da where d = (x,y) 0<x<4 and 0<y< X


1
Expert's answer
2021-11-01T11:43:10-0400

"∬ \frac{y}{x^2+1}da 
\\=∬ \frac{y}{x^2+1} dydx
\\=\int_{0}^{4}(\int_{0}^{x}\frac{y}{x^2+1}dy)dx
\\=\int_{0}^{4} \frac{1}{2}(\frac{y^2}{x^2+1})^{x}_{0}dx
\\=\frac{1}{2}\int_{0}^{4} \frac{x^2}{x^2+1}dx
\\=\frac{1}{2}\int_{0}^{4} \frac{x^2+1-1}{x^2+1}dx
\\=\frac{1}{2}[\int_{0}^{4} (1-\frac{1}{x^2+1})]dx
\\=\frac{1}{2}[\int_{0}^{4} (dx-\frac{1}{x^2+1}dx)]
\\=\frac{1}{2}[ x-tan^{-1}x]_{0}^{4}
\\=\frac{1}{2}[ 4-tan^{-1}4-0]
\\=2-\frac{1}{2}tan^{-1}4"